# A Survey on Self-play Methods in Reinforcement Learning

### v1

@@ Abstract
Self-play, characterized by agents’ interactions with copies or past versions of itself, has recently gained prominence in reinforcement learning. This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts. Then it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios. Finally, the survey highlights open challenges and future research directions in self-play. This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
Index Terms:
Self-play, reinforcement learning, game theory, multi-agent
I Introduction
Reinforcement learning (RL) represents a significant paradigm [1] within machine learning, concerned with the optimization of decision-making processes through interaction with an environment. It’s fundamentally modeled using a Markov decision process (MDP), a mathematical framework that describes an environment in terms of states, actions, transitions, and rewards. Within an MDP, agents operate by observing states, executing actions according to defined policies, receiving subsequent rewards, and transitioning to subsequent states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators [2]. This fusion of deep learning with RL has been instrumental in handling high-dimensional state spaces, contributing to breakthroughs in various complex tasks.
Moreover, the transition from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [3, 4, 5]. In MARL, the interdependence of agents’ actions introduces significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. By addressing issues such as non-stationarity and coordination, self-play offers an approach where an agent interacts with copies or past versions of itself [6, 7]. This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [8, 9, 10, 11], chess [10, 11], poker [12, 13], and video games [14, 15]. In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[8, 10].
Although some research takes a broad perspective through empirical game-theoretic analysis (EGTA) [16], it is important to note that there are relatively few comprehensive surveys focusing exclusively on self-play. Among these, some studies address the theoretical safety of self-play [17], while others develop an algorithmic framework for self-play that unfortunately does not accommodate the Policy-Space Response Oracle (PSRO) series of algorithms [18]. Furthermore, another study concentrates exclusively on PSRO [19]. Although these varied studies are valuable, they do not offer a perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to bridge this gap.
The survey is organized as follows. Sec. II introduces the background of self-play, including the RL framework and basic game theory concepts. Sec. III proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework, clarifying the self-play landscape. In Sec. IV, a comprehensive analysis is performed to illustrate how self-play is applied in various scenarios. Sec. V describes open problems in self-play and explores future research directions. Finally, Sec. VI concludes the survey on self-play.
II Preliminaries
In this section, we first introduce the framework of RL. Next, we present the basic game theory concepts and the typical evaluation metrics used in self-play.
II-A RL Framework
@@
In the RL context, agents interact with the environment based on the subsequent protocol: At each discrete time step $t$, every agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ is the parameters. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment undergoes a transition from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to maximize the expected discount accumulated rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
II-B Game Theory Concepts
II-B1 (Im)Perfect Information and (In)Complete Information
In a game characterized by perfect information, only one player moves at a time. Each player has a comprehensive understanding of the current game state, the full history of moves that have been made, and all potential future developments. If these conditions are not met, the game is considered to have imperfect information [25, 26]. In a game of incomplete information, there exists at least one player who is unaware of the payoff of another player; otherwise, it is a game of complete information [27].
For instance, Go is a game of both perfect and complete information. Players have full awareness of the entire game structure, including all possible moves, and they can see every move made by their opponent as they take turns to act (perfect information). Furthermore, if the outcomes are considered binary, such as win or loss, the payoff for the players is known to both sides (complete information).
II-B2 Normal-Form and Extensive-Form
@@
The normal-form and extensive-form are two different ways of representing games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ is pure strategy space of all players. A vector $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A pure strategy defines a specific and deterministic action for a player in a game, while a mixed strategy designates a probability distribution over the set of pure strategies, allowing randomized actions. A mixed strategy for the player $i$ is a probability distribution $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ is a probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$. If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game. If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game. If finite players (especially two players) are involved and each player has a finite set of strategies, a game in the normal-form can be directly depicted in a matrix.
@@
Specifically, in two-player zero-sum symmetric normal-form games, the pure strategy space for both player 1 and player 2 is identical, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. As the utility function $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, for simplicity, we can use only one utility function $u$ such that $\forall\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. The evaluation matrix captures the outcomes of the game by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
@@
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is chance and can be regarded as a special agent. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
@@
The Prisoner’s Dilemma serves as a classic example to illustrate various concepts in game theory. In a modified version of the dilemma, the outcomes are as follows:
@@
The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must decide simultaneously whether to confess or lie, without knowledge of the other’s choice. Games played in this manner are referred to as static games. The normal-form representation is particularly suitable for static games, as it captures the simultaneous nature of decision-making (as depicted in Fig. 1a). Another variant is the sequential Prisoner’s Dilemma, where the second player makes their decision with knowledge of the first player’s action. Games of this nature are called dynamic games. The extensive-form representation is well-suited for dynamic games, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig. 1d).
Furthermore, a normal-form representation can be transformed into an extensive-form representation. For instance, the transformation from Fig. 1a to Fig. 1b involves representing the simultaneous decisions in a tree structure. In this extensive representation, the dotted lines indicate that certain statuses belong to the same information set. Conversely, an extensive-form representation can also be transformed into a normal-form representation. For example, the transformation from Fig. 1d to Fig. 1c involves condensing the sequential decisions into a matrix format. In this case, the pure strategy space for player 2 is $2\times 2=4$, reflecting the need for player 2 to have a strategy for each of player 1’s possible actions (confess or lie). For instance, the notation for one of player 2’s pure strategies $C\rightarrow C,L\rightarrow C$ in Fig. 1c indicates that player 2 chooses to confess when he knows player 1 has chosen to confess and player 2 also chooses to confess when he knows player 1 has chosen to lie. Other symbols in Fig. 1c follow the same logic.
(a) Matrix Representation of simultaneous Prisoner’s Dilemma in normal-form.
@@
(b) Game Tree Representation of simultaneous Prisoner’s Dilemma in extensive-form.
(c) Matrix Representation of sequential Prisoner’s Dilemma in normal-form.
(d) Game Tree Representation of sequential Prisoner’s Dilemma in extensive-form.
Figure 1: The example of Prisoner’s Dilemma.
Compared to normal-form games, extensive-form games introduce sequential decision-making, adding complexity to the game structure. Extensive-form games also have a close relationship with MGs. In MGs with simultaneous moves, agents’ actions are unknown to each other, creating various histories that are condensed into a single information set. The game’s utility is the sum of rewards discounted over time [28]. Beyond normal-form and extensive-form games, as well as MGs, the analysis of complex Markov or extensive-form games often employs a higher-level abstraction: the meta-game. The meta-game facilitates the exploration of policy learning within these games, focusing not on isolated actions but on broader strategies that arise from the game’s dynamics. In this advanced normal-form context, the policy population consists of the strategies currently employed by players. Meta-strategies are mixed strategies that assign probabilities over the policy population in the meta-game.
II-B3 Transitive Game and Non-transitive Game

### v2

@@ Abstract
Self-play, characterized by agents’ interactions with copies or past versions of themselves, has recently gained prominence in reinforcement learning (RL). This paper first clarifies the preliminaries of self-play, including the multi-agent reinforcement learning framework and basic game theory concepts. Then, it provides a unified framework and classifies existing self-play algorithms within this framework. Moreover, the paper bridges the gap between the algorithms and their practical implications by illustrating the role of self-play in different scenarios. Finally, the survey highlights open challenges and future research directions in self-play. This paper is an essential guide map for understanding the multifaceted landscape of self-play in RL.
keywords:
@@
1 Introduction
Reinforcement learning (RL) represents a significant paradigm [120] within machine learning (ML), focused on optimizing decision processes through interaction with an environment. In RL, the environment is modeled as a Markov decision process (MDP), where agents observe states, take actions, receive rewards, and cause transitions to new states. The primary goal of RL algorithms is to derive the optimal policy that yields the maximum expected accumulated reward over time. Deep RL extends traditional RL by employing deep neural networks as function approximators to handle high-dimensional state spaces, contributing to breakthroughs in various complex tasks [94].
Moreover, transitioning from single-agent to multi-agent reinforcement learning (MARL) introduces complex dynamics [107, 85, 145]. In MARL, the interdependence of agents’ actions presents significant challenges, as the environment appears non-stationary to each agent. The main issues in MARL are coordination, communication, and equilibrium selection, particularly in competitive scenarios. These challenges often lead to difficulties in achieving convergence, maintaining stability, and efficiently exploring the solution space.
With the help of game theory, a mathematical framework that models the interactions between multiple decision-makers, self-play emerges as an elegant solution to some inherent challenges in MARL. Self-play provides an approach where an agent interacts with copies or past versions of itself [108, 5]. This method promises a more stable and manageable learning process. The capabilities of self-play extend to a wide range of scenarios, including its high-profile applications in Go [115, 117, 116, 112], chess [116, 112], poker [95, 50], and video games [7, 129]. In these scenarios, it has developed strategies that surpass human expertise. Although the application of self-play is extensive and promising, it is accompanied by limitations, such as the potential convergence to suboptimal strategies and significant computational requirements[115, 116].
Although some research adopts a broad perspective through empirical game-theoretic analysis (EGTA)[134], relatively few comprehensive surveys focus solely on self-play algorithms. Among them, a study develops an algorithmic framework for self-play, but it does not incorporate the Policy-Space Response Oracle (PSRO) series of algorithms [54], while a separate study focuses exclusively on PSRO [8], without considering other self-play algorithms. Another study examines the theoretical safety of self-play [34]. While these studies are valuable, they do not provide a comprehensive perspective that fully captures the breadth and depth of self-play. Therefore, this survey aims to fill this gap.
The survey is organized as follows. Sec. 2 introduces the background of self-play, including the RL framework and game theory concepts. Sec. 3 proposes a unified framework and then categorizes existing self-play algorithms into four categories based on this framework. In Sec. 4, a comprehensive analysis illustrates how self-play is applied in various scenarios. Sec. 5 describes open problems in self-play and explores future research directions. Finally, Sec. 6 concludes the survey on self-play. A more detailed overview of our survey is depicted in Fig. 1, which further illustrates the relationships among the subsections.
Figure 1: Overview of our survey.
2 Preliminaries
This section first introduces the framework of RL, where an agent learns to make decisions by interacting with an environment to maximize cumulative rewards. We then present basic game theory concepts, which analyze strategic interactions between rational decision-makers and provide tools to study outcomes based on players’ strategies.
2.1 RL Framework
In RL, the environment is modeled as an MDP, where the Markovian assumption states that the environment’s evolution is fully determined by its current state, eliminating the need to consider past states. An agent interacts with the environment by taking actions, which result in different states and associated rewards. MDPs can be extended to multi-agent settings, known as Markov games (MGs) [79] or stochastic games [113]. We focus on the most general form: partially observable Markov games (POMGs), where multiple agents interact with the environment, and each agent receives only individual observations instead of the full state, meaning they have limited access to the environment’s state.
A POMG $\mathcal{G}$ is defined by $\mathcal{G}=(\mathcal{N},\mathcal{S},\mathcal{A},\mathcal{O},\mathcal{P},\mathcal{R},\gamma,\rho)$. $\mathcal{N}=\{1,\cdots,n\}$ denotes n agents. $\mathcal{S}$ is the state space. $\mathcal{A}=\prod_{i=1}^{n}\mathcal{A}_{i}$ is the product of the action space of each agent. Similarly, $\mathcal{O}=\prod_{i=1}^{n}\mathcal{O}_{i}$ is the product of the observation space of each agent. $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ denotes the transition probability from one state to another given the actions of each agent. =$\mathcal{R}=\{\mathcal{R}_{1},\cdots,\mathcal{R}_{n}\}$, where $\mathcal{R}_{i}:\mathcal{S}\times\mathcal{A}_{i}\rightarrow\mathbb{R}$ denotes the reward function of agent $i$. $\gamma\in[0,1]$ is the discount factor. $\rho:\mathcal{S}\rightarrow[0,1]$ describes initial state distribution. If it is a fully cooperative setting, agents can share the same reward function [145, 106, 39, 68]. When $n=1,\mathcal{O}_{i}=\mathcal{S}$, the POMG returns to the MDP.
@@ 2.1 RL Framework
More concretely, in RL, agents interact with the environment through the following procedure: At each discrete time step $t$, each agent $i$ receives an observation $o_{i,t}$ from the environment and selects an action based on a stochastic policy $\pi_{\theta_{i}}:\mathcal{O}_{i}\times\mathcal{A}_{i}\rightarrow[0,1]$, where $\theta_{i}$ represents the parameters. After receiving the joint actions $\mathbf{a}_{t}=(a_{1,t},\cdots,a_{n,t})$, the environment transitions from the current state $s_{t}$ to a subsequent state $s_{t+1}$ according to the transition function $\mathcal{P}$ and sends a reward $r_{i,t+1}$ to every agent $i$. The ultimate goal of agent $i$ is to maximize the expected discounted cumulative rewards: $\mathbb{E}_{\pi_{\theta_{i}}}[\sum_{t=0}^{\infty}\gamma^{t}r_{i,t}]$.
2.2 Game Theory Concepts
2.2.1 Perfect Information and Imperfect Information
In a game with perfect information, only one player moves at a time. Each player comprehensively understands the current game state, the complete history of moves made, and the potential future outcomes. If these conditions are not met, the game is considered to have imperfect information [84, 99]. An example of a game with perfect information is Go. In Go, both players have full access to the game state at all times, including the positions of all stones and the identity of the next player. In contrast, Texas Hold’em poker is a game with imperfect information. Players cannot see their opponents’ private cards, creating uncertainty and requiring decisions based on probabilistic reasoning, bluffing, and interpreting others’ actions.
2.2.2 Normal-Form and Extensive-Form
@@
The normal form and extensive form are two distinct representations of games in game theory. If a game $\mathcal{G}$ is represented in the normal-form, it can be expressed by $\mathcal{G}=(\mathcal{N},\mathbf{\Pi},\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes the players. The set $\mathbf{\Pi}=\Pi_{1}\times\cdots\times\Pi_{n}$ represents the pure strategy space for all players. A pure strategy specifies a particular and deterministic action for a player in the game. A vector $\bm{\pi}=(\pi_{1},\dots,\pi_{n})\in\mathbf{\Pi}$ is called a strategy profile. A mixed strategy assigns a probability distribution over the set of pure strategies. For player $i$, a mixed strategy is represented by $\sigma_{i}\in\Delta(\Pi_{i})$, where $\Delta$ denotes the probability simplex. $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:\mathbf{\Pi}\rightarrow\mathbb{R}$, is a utility function that assigns a real-valued payoff to each player $i$. If $\forall\bm{\pi}\in\mathbf{\Pi},\sum_{i}u_{i}(\bm{\pi})=0$, the game is a zero-sum game, otherwise it is a general-sum game. If $\Pi_{1}=\cdots=\Pi_{n}$ and the payoffs are invariant under any permutation of the players’ strategies, the game is a symmetric game. If a finite set of players (especially two players) are involved and each player has a finite set of strategies, a normal-form game can be directly depicted in a matrix.
@@
In two-player zero-sum symmetric normal-form games, both players share the same pure strategy space, denoted by $\Pi$, such that $\Pi=\Pi_{1}=\Pi_{2}$. Since the utility function satisfies $u_{1}(\pi_{i},\pi_{j})=-u_{2}(\pi_{i},\pi_{j})$, we can simplify the utility to a single function $u$, where for $\pi_{i},\pi_{j}\in\Pi$, if $\pi_{i}$ beats $\pi_{j}$, then $u(\pi_{i},\pi_{j})=-u(\pi_{j},\pi_{i})>0$. The evaluation matrix captures the game outcomes by detailing the results of different strategies when they are played against each other: $A_{\Pi}=\{u(\pi_{i},\pi_{j}):\pi_{i},\pi_{j}\in\Pi\times\Pi\}$.
@@
If a game is represented in the extensive-form, it is expressed sequentially, illustrating the sequence of moves, choices made by the players, and the information available to each player during decision-making. Typically, a game in the extensive-form is represented by a game tree. This tree demonstrates the sequential and potentially conditional nature of decisions. Moreover, if player $i$ has perfect recall, it means that player $i$ remembers which action they have taken in the past. A game $\mathcal{G}$ represented in the extensive-form can be expressed by $\mathcal{G}=(\mathcal{N}\cup\{c\},H,Z,P,\mathcal{I},A,\mathbf{u})$. $\mathcal{N}=\{1,2,\cdots,n\}$ denotes a set of players. $c$ is chance node and can be regarded as a special player. $H$ represents a set of possible histories and $Z\subseteq H$ is a set of terminal histories. Order of moves is represented by a function $P(h)\in\mathcal{N}\cup\{c\}$ to indicate which player is to move, where $h\in H$. $\mathcal{I}$ denotes information set partitions and $\mathcal{I}_{i}$ denotes the information set partitions for player $i$. This implies that in an imperfect information game, if player $i$ reaches a history $h\in I_{i}$, where $I_{i}\in\mathcal{I}_{i}$ is a specific information set, player $i$ cannot distinguish which particular history $h\in I_{i}$ it is encountering. Action space is represented by $A(h)$ for a non-terminal history $h\in H$. For all non-terminal histories $h$ within an information set $I_{i}$, the available actions are the same; otherwise, they are distinguishable. Therefore, we use $A(I_{i})$ to represent the available actions for the information set $I_{i}$. Utility functions is denoted by $\mathbf{u}=(u_{1},\cdots,u_{n})$, where $u_{i}:Z\rightarrow\mathbb{R}$. Together, these components define the structure and dynamics of an extensive-form game. Moreover, A strategy profile in an extensive-form game can be expressed by $\bm{\pi}=(\pi_{1},\cdots,\pi_{n})$, where $\pi_{i}$ maps each $I_{i}\in\mathcal{I}_{i}$ to a probability distribution over $A(I_{i})$. A subgame of an extensive-form game is a portion of the game that starts from a single initial node, includes all successors of any node within the subgame, and contains all nodes in the same information set as any node in the subgame.
@@
The Prisoner’s Dilemma is a classic example in game theory. The game outcomes are as follows:
@@
The classic scenario is known as the simultaneous Prisoner’s Dilemma, where two players must simultaneously decide whether to confess or lie without knowing the other’s choice. Games played in this manner are referred to as static games. The normal-form representation is suitable for them, as it captures the simultaneous nature of decision-making (as depicted in Fig. 2a). Another variant is the sequential Prisoner’s Dilemma, where the second player decides with knowledge of the first player’s action. Games of this nature are called dynamic games. The extensive-form representation is well-suited for them, as it can clearly illustrate the sequence of moves and the information available to each player at each decision point (as shown in Fig. 2b).
@@
Figure 2: The example of Prisoner’s Dilemma. (a) Matrix representation of simultaneous Prisoner’s Dilemma in normal-form. (b) Game tree representation of sequential Prisoner’s Dilemma in extensive-form.
Beyond normal-form and extensive-form games, the analysis of complex games often involves a higher-level abstraction: the meta-game. The meta-game can be viewed as an advanced form of a normal-form game, but it focuses on the selection of policies (also called strategies) rather than isolated actions. Each player chooses from a set of policies, rather than a set of actions. This set of policies is referred to as the policy population. Meta-strategies are mixed strategies that assign probabilities to the policies within the policy population in the meta-game.
2.2.3 Transitive Game and Non-transitive Game